The lifetime of a machine is continuous on the interval (0; 40) with probability density
function f, where f(t) is proportional to (t+10)^-2, and t is the lifetime in years. Find
the proportionality constant that makes the f(t) a probability function.
1
Expert's answer
2020-06-01T19:18:30-0400

We have, the lifetime of a machine is continuous on the interval (0, 40) with probability density

function f, where

f(t) "\propto" (t+10)-2, t is the lifetime in years


Let, k be the proportionality constant. Then,


f(t) = k(t+10)-2, 0 < t < 40

= 0, otherwise


Now in order to f(t) become a probability function


"\int_0^{40} f(t)dt" = 1


i.e. "\int_0^{40} k(t+10)^{-2}dt" = 1


i.e. "k\int_0^{40} \frac{dt}{(t+10)^2}" = 1


Let, t + 10 = u

"\implies"dt = du

When t = 0, u = 10 and when t = 40, u = 50


"\therefore k\int_{10}^{50} \frac{du}{u^2}" = 1


i.e. "-k[\frac{1}{u}]_{10}^{50}" = 1


i.e. "-k(\frac{1}{50}-\frac{1}{10})" = 1


i.e. "\frac{4k}{50}" = 1


i.e. k = 50/4 = 25/2


Answer: The proportionality constant that makes f(t) a probability function is 25/2.